Primes and Congruences ch:primes

Prime Factorization Prime Numbers We call positive whole numbers the natural numbers and denote them by . Thus

= {1, 2, 3, 4, …}.

We call all the whole numbers, both positive and negative, the integers, and write

= {…, −2, −1, 0, 1, 2, …}.

They are denoted by because the German word for the integers is “Zahlen” (and 19th century German number theorists rocked).

definition If a, b ∈ then “a divides b” if ac = b for some c ∈.